OpenStudy (anonymous):

I need help Writing Equations Using Point-Slope Form... (3,5) with a slope if -3? Please help me

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - y1 = -3(x - x1) ... plug in the given slope m = -3 y - 5 = -3(x - 3) ... plug in x1 = 3 and y1 = 5, this is from the given point (3,5)